# Practitioners’ Perspectives on a Differential Privacy Deployment Registry

### v1

@@ Abstract
Differential privacy (DP)—a principled approach to producing statistical data products (e.g., summary statistics, machine learning models) with strong, mathematically provable privacy guarantees for the individuals in the underlying dataset—has seen substantial adoption in practice over the past decade. Applying DP requires making several implementation decisions, each with significant impacts on data privacy and/or utility. Hence, to promote shared learning and accountability around DP deployments, Dwork, Kohli, and Mulligan [1] proposed a public-facing repository (“registry”) of DP deployments. The DP community has recently started to work toward realizing this vision. We contribute to this effort by (1) developing a holistic, hierarchical schema to describe any given DP deployment and (2) designing and implementing an interactive interface to act as a registry where practitioners can access information about past DP deployments. We (3) populate our interface with 21 real-world DP deployments and (4) conduct an exploratory user study with DP practitioners ($n=16$) to understand how they would use the registry, as well as what challenges and opportunities they foresee around its adoption. We find that participants were enthusiastic about the registry as a valuable resource for evaluating prior deployments and making future deployments. They also identified several opportunities for the registry, including that it can become a “hub” for the community and support broader communication around DP (e.g., to legal teams). At the same time, they identified challenges around the registry gaining adoption, including the effort and risk involved with making implementation choices public and moderating the quality of entries. Based on our findings, we offer recommendations for encouraging adoption and increasing the registry’s value not only to DP practitioners, but also to policymakers, data users, and data subjects.
September 16, 2025 (Draft)
@@ 1 Introduction
Differential privacy (DP) [2] was first introduced in 2006 as a theoretical standard of privacy for computing statistical releases from datasets with sensitive information about individuals. In the nearly two decades since, researchers and practitioners have undertaken substantial work to bring DP to practical fruition. To date, DP has been deployed to produce data products ranging from summary statistics to machine learning models, by organizations such as the U.S. Census Bureau [3], Google [4], Apple [5], and the Wikimedia Foundation [6]. A recent survey found at least 42 deployments reported in papers or blog posts [7].
Deploying DP involves making a range of implementation choices, each with potential to significantly impact data privacy and/or utility. Under DP, statistical noise is injected into analyses to obscure individual-level contributions while preserving overall signal in the data. The “privacy loss parameter” $\epsilon$ often receives much attention for calibrating how much noise is added—and therefore controlling the privacy-accuracy tradeoff—but it is important to note that there are several other decisions (e.g., whether the deployment model is central, local, etc.; the choice of privacy unit; etc.) that are consequential to resulting guarantees, as outlined by a recent set of guidelines on evaluating DP guarantees [8].
@@
To promote shared learning and overcome these challenges, Dwork, Kohli, and Mulligan [1] proposed the idea of a public-facing repository of DP deployments. They argued that it should contain both technical implementation decisions as well as other sociotechnical factors, such as rationale for said decisions and the processes by which they were made. Since their proposal, appetite for a registry has grown in the privacy community, but the original vision has not yet been achieved. Furthermore, if a registry were to exist, it would raise several challenges around adoption and governance, which are difficult to assess without a working registry.
Thus, the community has started to take steps toward building a registry and realizing the vision of Dwork, Kohli, and Mulligan [1]. First, in 2021, Damien Desfontaines [9] created a blog post that lists several real-world deployments of DP and is periodically updated by Desfontaines with new deployments. Though the post has become a fairly well-known, trusted resource in the DP community, its focus is primarily on summarizing deployments at a high level and listing privacy parameters, omitting several factors that are crucial for fully understanding a DP deployment. Furthermore, the blog post medium does not scale to larger numbers of deployments and relies on a single maintainer rather than being community sourced.
More recently, in spring 2024, a team at the privacy start-up Oblivious started to work toward a more systematic, community-driven registry to be further developed in collaboration with the OpenDP open-source software project, of which we are a part. They created an initial prototype [10] that was also in a blog post format, but with a table of deployments and educational content about DP. The table captured a basic set of attributes of a DP deployment (privacy parameters, privacy unit, deployment model, etc.) in a non-hierarchical structure. Our work builds on and contributes to this effort.
@@ 1.1 Our contributions
In our work, we develop a newer and richer prototype of a DP deployment registry geared toward practitioners. Specifically:
  • We propose a schema for describing DP deployments. It builds on attributes from the registry initiated by Oblivious, and is based on existing literature that aims to synthesize DP concepts [8, 11, 12, 1, 13], our knowledge as DP researchers, and iterative feedback from DP experts.
@@
Much of these contributions have been incorporated back into the registry prototype started by Oblivious, which is being launched to accompany a proposal for the National Institute of Standards and Technology (NIST) to host the registry [14]. Conversely, the concurrent work on the prototype started by Oblivious has influenced our work in several ways.
Next, we conduct a user study ($n=16$) with DP practitioners, using our prototype registry as a technology probe [15] to explore how they might use the registry in practice and what challenges or opportunities they foresee around its adoption. In particular, we ask the following research questions (RQs):
@@
Summary of findings. We find that interacting with the registry did not always change participants’ overall perceptions of DP, but consistently supported them in discovering low-level details and asking new questions about prior deployments. Furthermore, we find that participants performed nuanced evaluations of prior deployments, focusing not only on technical implementation choices, but also on overall societal relevance and transparency of the deployment. When tasked with making a future deployment using the registry, participants based their implementation choices off similar deployments—but determined similarity in different ways. Participants approached trends across deployments with curiosity, but healthy skepticism when they did not align with their prior beliefs. Finally, participants identified several opportunities the registry could open up, including becoming a “hub” for the community and supporting broader communication around DP (e.g., to legal teams). At the same time, they identified challenges related to the effort and risk involved with adding new deployments to the registry and moderating quality of entries.
We end with recommendations based on our findings for encouraging adoption and increasing the registry’s value to parties other than practitioners (e.g., policymakers, data users, data subjects).
@@ 1.2 Other related work
There have been multiple repositories focused on privacy enhancing technologies (PETs) broadly. The PETs use case repository by the Centre for Data Ethics and Innovation [16] contains case studies of real-world or proposed deployments of PETs. DP is included, but not the focus. Furthermore, the repository focuses on describing deployments at a high level and instead providing links to external write-ups. Similarly, the Repository of PETs Use Cases by the United Kingdom government [17] contains real-world uses of PETs and provides high-level descriptions with links. Last, the Future of Privacy Forum’s Repository for PETs [18] contains links to case studies as well as links to documents related to regulatory activity and other topics.
There have also been efforts to increase transparency around privacy practices generally. Nutrition labels [19] are an approach to documenting and communicating organizations’ data-sharing practices with data subjects; this approach has been extended to a DP context to convey information flows permitted under different deployment models [20]. Privacy nutrition labels, however, are aimed at data subjects who do not necessarily have a technical background, and they do not holistically capture aspects of DP deployments in particular.
Concurrent to our work, Khavkin and Toch [7] conducted a systematic literature review of DP deployments across academic, government, and commercial contexts. They propose a taxonomy of “key DP configuration factors” and study trends across the deployments identified in their search. Our schema includes these factors, as well as multiple additional details.
Finally, less immediately related to our work are proposals or methods of documenting machine learning models and datasets. Model cards [21] were proposed as an approach to documenting machine learning models, including their intended uses and performance across different groups of people, while datasheets [22] aim to document intended uses and other details of datasets. These two approaches do not focus specifically on privacy aspects of models or datasets, and do not provide a structure for how DP deployments in particular should be reported. However, we draw inspiration from model cards’ format, and create an adaptation to the DP context.
@@ 2 Differential privacy deployment card
To create a registry, we first develop a schema for any given DP deployment. The schema enables deployments to be represented in a standardized form and facilitates comparisons across them. We developed the schema, shown below as a template for a “deployment card” (see Figure 1), by triangulating between and synthesizing recommendations from multiple sources: the original proposal for a registry [1], which lists attributes of deployments that the registry should describe, NIST’s guidelines for evaluating DP guarantees [8], an applied guidebook for DP practitioners [13], and two approaches to describing DP flavors more generally (OpenDP’s programming framework [11] and [12]). Our schema includes a wider range of attributes than what was originally proposed by Dwork, Kohli, and Mulligan [1] and is hierarchical, offering additional organization and structure. We used an iterative process to develop the schema, discussing additions within our research team and seeking feedback from three external DP experts.
@@ 2.1 Deployment card
In this section, we introduce sections of the deployment card and describe their contents. Furthermore, we propose that deployment cards can be filled out according to different “transparency tiers” (Tier 1, 2, or 3). Each increasing tier requires more information to be disclosed about the deployment, allowing for a tradeoff between ease of registration and greater transparency. Below, we also describe which sections are recommended for each tier. Last, we use the term “underlying data” to refer to the data used to produce the data product; in other words, the data being protected.
Differential privacy deployment card Data product Name Data curator Description Publication date Region Sector Flavor Flavor name (e.g., pure, approximate) Data domain Unprotected quantities Privacy loss Privacy unit Privacy parameters Deployment model Model name (e.g., central, local) Trust assumptions Release type (one release vs. many releases) Data source (dynamic vs. static) Access type (interactive vs. non-interactive) Accounting Composition Post-processing Implementation Pre-processing, exploratory data analysis, hyperparameter tuning Mechanisms Justification More information & sources Links to papers, blogposts, or other resources describing the deployment Link to the data product, if available Additional notes about the deployment
@@ 2.1.1 Data product (basic information)
Region. The region in which the underlying data were collected. In other words, location in which the data subjects are based.
Sector. The sector of the data product, such as technology, healthcare, education, government, or energy.
@@ 2.1.2 Flavor
This section describes the flavor of DP employed as well as related attributes. We suggest that Tier 2 entries specify the flavor name, while the data domain and unprotected quantities are also recommended for Tier 3 entries.
Flavor name. The name of the DP flavor (e.g., pure DP [2], approximate DP [23], zero-concentrated DP [24], Rényi DP [25]).
Data domain. Underlying datasets “eligible for privacy protection”; in other words, “the actual, potential, or counterfactual datasets that are to be protected” under DP [12].
Unprotected quantities. Any quantities in the data product that are unprotected by DP (e.g., statistics computed over a dataset that are published in the clear, i.e., without DP noise). These quantities are sometimes also referred to as “invariants.”
@@ 2.1.3 Privacy loss
This section describes the strength of DP guarantees. Both the privacy unit and privacy parameters are required for Tier 2 entries, but we suggest providing a more detailed explanation of the privacy unit for Tier 3 entries, as described below.
Privacy unit. The entity being protected. For example, the privacy unit might be a person or household, or a single event (“event-level” privacy) or the set of all events associated with a particular user in one day (“user-day” privacy). In other words, the privacy unit “characterizes what [is being] protected” [13]. More formally, the privacy unit is the entity whose data changes under adjacent datasets [12]. Tier 3 entries should go beyond stating a high-level description of the privacy unit and include a precise specification of what constitutes adjacent datasets.
Privacy parameters. Values for parameters like $\epsilon$, $\delta$, and $\rho$; the specific parameters that apply to a given deployment will vary according to the DP flavor used. Privacy parameters have been described as expressing the “intensity of protection” [12].
@@ 2.1.4 Deployment model
Model name. The name of the DP deployment model (e.g., central, local [2, 26]). Each deployment model implies different types or levels of trust in the data curator and other relevant actors. Each model also implies different points in the deployment pipeline at which DP noise is added. For instance, in the central model, a trusted data curator collects or is otherwise permitted to access data as collected from data subjects. The curator is responsible for applying DP to create a privacy-protected data product that can be shared with other parties who are not permitted to see the underlying data. On the other hand, in the local model [2, 26], the curator is not permitted to access the underlying data. Instead, DP is applied to data at the individual level, before it is sent from each data subject to the curator. The curator may then be responsible for combining data subjects’ privacy-protected data in a meaningful way.
@@
Release type. Whether the data product is published once (“one release”) or many times (“many releases”). For deployments that involve one release, Tier 3 entries should state plans for future uses or publications of the underlying data, if any. For deployments that involve many releases, Tier 3 entries should include a description of the refreshment timeframe (the amount of time after which the privacy loss budget resets), how privacy loss is managed over time, and whether a fixed amount of privacy loss is allowed before the underlying data are no longer queried.
Data source. Refers to the static or dynamic nature of the underlying data. If the underlying data are dynamic, new underlying data come in over time and new releases are periodically made (so this is always coupled with “Release type” being “many releases”). This is also known as the setting of DP under continual observation. On the other hand, if the underlying data are static, there is a single, unchanging dataset on which one or many releases are made.
Access type. Refers to how potential users of the data product may access it; this attribute is either “interactive” or “non-interactive.” Under interactive deployments, people with permission, like data analysts, can interactively query the underlying data under DP, and receive privacy-protected query estimates. Tier 3 entries that are interactive should also describe how the privacy loss budget is apportioned to and across analysts, and whether non-collusion between analysts is assumed. Under non-interactive deployments, people cannot interactively query the underlying data. Instead, they must interact with the published data product as is.

### v2

@@ Abstract
Differential privacy (DP)—a principled approach to producing statistical data products (e.g., summary statistics, machine learning models) with strong, mathematically provable privacy guarantees for the individuals in the underlying dataset—has seen substantial adoption in practice over the past decade. Applying DP requires making several implementation decisions, each with significant impacts on data privacy and/or utility. Hence, to promote shared learning and accountability around DP deployments, Dwork, Kohli, and Mulligan dwork2019differential proposed a public-facing repository (“registry”) of DP deployments. The DP community has recently started to work toward realizing this vision. We contribute to this effort by (1) developing a holistic, hierarchical schema to describe any given DP deployment and (2) designing and implementing an interactive interface to act as a registry where practitioners can access information about past DP deployments. We (3) populate our interface with 21 real-world DP deployments and (4) conduct an exploratory user study with DP practitioners ($n=16$) to understand how they would use the registry, as well as what challenges and opportunities they foresee around its adoption. We find that participants were enthusiastic about the registry as a valuable resource for evaluating prior deployments and making future deployments. They also identified several opportunities for the registry, including that it can become a “hub” for the community and support broader communication around DP (e.g., to legal teams). At the same time, they identified challenges around the registry gaining adoption, including the effort and risk involved with making implementation choices public and moderating the quality of entries. Based on our findings, we offer recommendations for encouraging adoption and increasing the registry’s value not only to DP practitioners, but also to policymakers, data users, and data subjects.
November 24, 2025
@@ 1 Introduction
Differential privacy (DP) dwork2006calibrating was first introduced in 2006 as a theoretical standard of privacy for computing statistical releases from datasets with sensitive information about individuals. In the nearly two decades since, researchers and practitioners have undertaken substantial work to bring DP to practical fruition. To date, DP has been deployed to produce data products ranging from summary statistics to machine learning models, by organizations such as the U.S. Census Bureau abowd2018us, Google aktay2020google, Apple apple2017, and the Wikimedia Foundation adeleye2023publishing. A recent survey found at least 42 deployments reported in papers or blog posts khavkin2025differential.
Deploying DP involves making a range of implementation choices, each with potential to significantly impact data privacy and/or utility. Under DP, statistical noise is injected into analyses to obscure individual-level contributions while preserving overall signal in the data. The “privacy loss parameter” $\epsilon$ often receives much attention for calibrating how much noise is added—and therefore controlling the privacy-accuracy tradeoff—but it is important to note that there are several other decisions (e.g., whether the deployment model is central, local, etc.; the choice of privacy unit; etc.) that are consequential to resulting guarantees, as captured by a recent set of guidelines on evaluating DP guarantees NIST.SP.800-226.
@@
To promote shared learning and overcome these challenges, Dwork, Kohli, and Mulligan dwork2019differential proposed the idea of a public-facing repository of DP deployments. They argued that it should contain both technical implementation decisions and sociotechnical factors, such as rationale for said decisions and the processes by which they were made. Appetite for a registry has grown in the privacy community, but the original vision has not yet been achieved. Furthermore, if a registry were to exist, it would raise several challenges around adoption and governance, which are difficult to assess without a working registry.
Thus, the community has started to take steps toward building a registry and realizing the vision of Dwork, Kohli, and Mulligan dwork2019differential. First, in 2021, Desfontaines desfontainesblog20211001 created a blog post that lists several real-world deployments of DP and is periodically updated by Desfontaines with new deployments. Though the post has become a fairly well-known, trusted resource in the DP community, its focus is primarily on summarizing deployments at a high level and listing privacy-loss parameters, omitting several factors that are crucial for fully understanding a DP deployment. Furthermore, the blog post medium does not scale to larger numbers of deployments and relies on a single maintainer rather than being community sourced.
More recently, in spring 2024, a team at the privacy start-up Oblivious (Berrios, Usynin, Fitzsimons original_registry) started to work toward a more systematic, community-driven registry to be further developed in collaboration with OpenDP, the open-source software project of which we are a part. They created an initial prototype that was also in a blog post format, but with a table of deployments and educational content about DP. The table captured a basic set of attributes of a DP deployment (privacy-loss parameters, privacy unit, deployment model, etc.) in a non-hierarchical structure. Our work builds on and contributes to this effort.
@@ 1.1 Our contributions
We develop a newer, richer prototype of a DP deployment registry geared toward practitioners. Specifically:
  • We propose a schema for describing DP deployments. It builds on attributes from the registry initiated by Oblivious, and is based on existing literature that aims to synthesize DP concepts NIST.SP.800-226, gaboardi2020programming, bailie2025topics, dwork2019differential, cowan2024hands, our knowledge as DP researchers, and iterative feedback from DP experts.
@@
Much of these contributions were incorporated back into the registry prototype started by Oblivious, which was launched in September 2025 to accompany a proposal for the U.S. National Institute of Standards and Technology (NIST) to host the registry NIST.IR.8588.2025. Conversely, the concurrent work on the prototype started by Oblivious has influenced our work in several ways.
Next, we conduct a user study ($n=16$) with DP practitioners, using our prototype registry as a technology probe hutchinson2003technology to explore how they might use the registry in practice and what challenges or opportunities they foresee around its adoption. In particular, we ask the following research questions (RQs):
@@
Summary of findings. We find that interacting with the registry did not always change participants’ overall perceptions of DP, but consistently supported them in discovering low-level details and asking new questions about prior deployments. Furthermore, we find that participants performed nuanced evaluations of prior deployments, focusing not only on technical implementation choices, but also on overall societal relevance and transparency of the deployment. When tasked with making a future deployment using the registry, participants based their implementation choices off similar deployments—but determined similarity in different ways. Participants approached trends across deployments with curiosity, but also skepticism when they did not align with their prior beliefs. Finally, participants identified several opportunities the registry could open up, including becoming a “hub” for the community and supporting broader communication around DP (e.g., to legal teams). At the same time, they identified challenges related to the effort and risk involved with adding new deployments to the registry and moderating quality of entries.
We end with recommendations based on our findings for encouraging adoption and increasing the registry’s value to parties beyond practitioners (e.g., policymakers, data users, data subjects).
@@ 1.2 Other related work
There have been multiple repositories focused on privacy enhancing technologies (PETs) broadly. The PETs Repository of Use Cases by the Centre for Data Ethics and Innovation cdeiukRepositoryCases contains case studies of real-world or proposed deployments of PETs. DP is included, but is not the focus. Furthermore, the repository focuses on describing deployments at a high level and providing links to external write-ups. Similarly, the Repository of PETs Use Cases by the U.K. government UK_PETS_Repository contains real-world uses of PETs and provides high-level descriptions with links. Last, the Future of Privacy Forum’s Repository for PETs fpfRepository contains links to case studies and documents related to regulatory activity and other topics.
There have also been efforts to increase transparency around privacy practices generally. Nutrition labels kelley2009nutrition are an approach to documenting and communicating organizations’ data-sharing practices typically with data subjects; this approach has been extended to a DP context to convey information flows permitted under different deployment models smart2025models. Concurrent to our work, Dibia, Lu, Bhattacharjee, Near, and Feng dibia2025we developed a DP privacy label for technical audiences; our schema contains the majority of information in their label and additional details.
Also concurrent to our work, Khavkin and Toch khavkin2025differential conducted a systematic literature review of DP deployments across academic, government, and commercial contexts. They propose a taxonomy of “key DP configuration factors” and study trends across deployments. Our schema includes these factors and multiple additional details.
Finally, less immediately related to our work are proposals or methods of documenting machine learning models and datasets. Model cards mitchell2019model were proposed as an approach to documenting machine learning models—including intended uses and performance across different groups of people—while datasheets gebru2021datasheets aim to document intended uses and other details of datasets. These two approaches do not focus specifically on privacy aspects of models or datasets, and do not provide a structure for how DP deployments should be reported. However, we draw inspiration from model cards’ format, and create an adaptation to the DP context.
@@ 2 Differential privacy deployment card
To create a registry, we first develop a schema for any given DP deployment. The schema enables deployments to be represented in a standardized form and facilitates comparisons across them. We developed the schema, summarized as a “deployment card” (Figure 1), by triangulating between and synthesizing recommendations from multiple sources: the original proposal for a registry by Dwork, Kohli, and Mulligan dwork2019differential, which lists attributes of deployments that the registry should describe; NIST’s guidelines for evaluating DP guarantees by Near, Darais, Lefkovitz and Howarth NIST.SP.800-226; an applied guidebook for DP practitioners by Cowan, Shoemate, and Pereira cowan2024hands; and two approaches to describing DP flavors more generally, specifically OpenDP’s programming framework by Gaboardi, Hay, and Vadhan  gaboardi2020programming and one by Bailie bailie2025topics. Our schema includes a wider range of attributes than what was originally proposed in Dwork, Kohli, and Mulligan dwork2019differential and is hierarchical, offering additional detail and structure. We used an iterative process to develop the schema, discussing additions within our research team and seeking feedback from three external DP experts.
@@ 2.1 Deployment card
Here, we introduce and describe sections of the deployment card. We propose that deployment cards can be filled out according to different “transparency tiers” (Tier 1, 2, or 3). Each increasing tier requires more information to be disclosed about the deployment, allowing for a tradeoff between ease of registration and greater transparency. Below, we describe which sections are recommended for each tier. Throughout, we use the term “underlying data” to refer to the data used to produce the data product (i.e., the data being protected). Finally, we present the deployment card as it was instantiated in our prototype to accurately contextualize our user study findings. Throughout this section, we also include footnotes with possible modifications or improvements for future integration into the deployment card and registry.
Differential privacy deployment card Data product Name Data curator Description Intended use Publication date Region Sector Flavor Flavor name (e.g., pure, approximate) Data domain Unprotected quantities Privacy loss Privacy unit Privacy-loss parameters Deployment model Model name (e.g., central, local) Trust assumptions Release type (one release vs. many releases) Data source (dynamic vs. static) Access type (interactive vs. non-interactive) Accounting Composition Post-processing Implementation Pre-processing, exploratory data analysis, hyperparameter tuning Mechanisms Justification More information & sources Links to papers, blogposts, or other resources describing the deployment Link to the data product, if available Additional notes about the deployment
@@ 2.1.1 Data product (basic information)
Region. The region in which the underlying data were collected (i.e., location in which the data subjects are based).
Sector. The sector of the data product, such as technology, healthcare, education, government, or energy. [^1] A future schema may also include an attribute for sector of the curator.
@@ 2.1.2 Flavor
This section describes the flavor of DP employed and related attributes. We suggest that Tier 2 entries specify the flavor name, while the data domain and unprotected quantities are also recommended for Tier 3 entries.
Flavor name. [^2] For precision, rename “Flavor name” to “privacy measure.” The name of the DP flavor (e.g., pure DP dwork2006calibrating, approximate DP dwork2006our, zero-concentrated DP bun2016concentrated, Rényi DP mironov2017renyi, Gaussian DP dong2022gaussian).
Data domain. Underlying datasets “eligible for privacy protection”, i.e., “the actual, potential, or counterfactual datasets that are to be protected” under DP bailie2025topics.
Unprotected quantities. Any quantities in the data product that are unprotected by DP (e.g., statistics computed over a dataset that are published without DP noise). These quantities are sometimes referred to as “invariants.”
@@ 2.1.3 Privacy loss
This section [^3] Instead of creating a separate section for privacy loss, an alternative is to group privacy unit with data domain (because the privacy unit is the entity whose data changes under adjacent datasets in the data domain bailie2025topics) and privacy-loss parameters with the privacy measure because the privacy measure defines the meaning of the privacy-loss parameters. describes the strength of DP guarantees. Both the privacy unit and privacy-loss parameters are required for Tier 2 entries, but we suggest providing a more detailed explanation of the privacy unit for Tier 3 entries, as described below.
Privacy unit. The entity being protected. For example, the privacy unit might be a person or household, or a single event (“event-level” privacy) or the set of all events associated with a particular user in one day (“user-day” privacy). The privacy unit “characterizes what [is being] protected” cowan2024hands. More formally, it is the entity whose data changes under adjacent datasets bailie2025topics; it is defined by the input metric and maximum input distance gaboardi2020programming. Tier 3 entries should also include a precise specification of what constitutes adjacent datasets.
Privacy-loss parameters. Values for parameters like $\epsilon$, $\delta$, and $\rho$. The specific parameters that apply to a given deployment will vary according to the DP flavor used. Privacy-loss parameters express the “intensity of protection” bailie2025topics.
@@ 2.1.4 Deployment model
Model name. The name of the DP deployment model (e.g., central dwork2006calibrating, local dwork2006calibrating, kasiviswanathan2011can, shuffle cheu2019distributed, bittau2017prochlo, federated beimel2008distributed). Each deployment model implies different types or levels of trust in the data curator and other relevant actors. Each model also implies different points in the deployment pipeline at which DP noise is added. For instance, in the central model, a trusted data curator collects or is otherwise permitted to access raw data from data subjects. The curator is responsible for applying DP to create a privacy-protected data product that can be shared with other parties who are not permitted to see the underlying data. In the local model dwork2006calibrating, kasiviswanathan2011can, the curator is not permitted to access the underlying data. Instead, DP is applied to data at the individual level, before it is sent from each data subject to the curator. The curator may then be responsible for combining data subjects’ privacy-protected data in a meaningful way.
@@
Release type. Whether the data product is published once (“one release”) or many times (“many releases”). For deployments that involve one release, Tier 3 entries should state plans for future uses or publications of the underlying data, if any. For deployments that involve many releases, Tier 3 entries should include a description of the refreshment timeframe (the amount of time after which the privacy loss budget resets), how privacy loss is managed over time, and whether a fixed amount of privacy loss is allowed before the underlying data are no longer queried. [^4] These details could alternatively appear in the accounting section, as they directly relate to how privacy loss is apportioned and analyzed.
Data source. Refers to the static or dynamic nature of the underlying data. If the underlying data are dynamic, new underlying data come in over time and new releases are periodically made (so this is always coupled with “Release type” being “many releases”). This is also known as the setting of DP under continual observation dwork2010differential. On the other hand, if the underlying data are static, there is a single, unchanging dataset on which one or many releases are made.
Access type. Refers to how potential users of the data product may access it; this attribute is either “interactive” or “non-interactive.” Under interactive deployments, people with permission, like data analysts, can interactively query the underlying data under DP, and receive privacy-protected query estimates. Tier 3 entries that are interactive should also describe how the privacy loss budget is apportioned to and across analysts, and whether non-collusion between analysts is assumed. [^5] Again, these details could appear in the accounting section. Under non-interactive deployments, people cannot interactively query the underlying data. Instead, they must interact with the published data product as is.